# Data-Adaptive Graph Framelets with Generalized Vanishing Moments for Graph Signal Processing

### v2

@@ IV-B Learning the High-pass Filters
Once the filters $\bm{A}^{[j,k]}$ are determined, in view of the condition in (17), we can obtain each $\bm{B}^{[j,k]}$ by finding a basis or a frame of the orthogonal complement of the rows of $\bm{A}^{[j,k]}$. For the former case, we can apply singular value decomposition (SVD). For the latter case, we can apply the method of constructing undecimated framelet systems in [44, Theorem 2] on a set of orthonormal eigenvectors and select three filters $\bm{\alpha},\bm{\beta}_{1},\bm{\beta}_{2}$ as in [46, Section 4.1]. Here, we choose not to describe the details for brevity and assume that such initial high-pass filters $\bm{B}^{[j,k]}$ are obtained either through the SVD technique or the undecimated framelet approach. We would like to emphasize that any method for constructing frames on the span of a set of orthonormal vectors is eligible.
@@
It is obvious that the high-pass filters $\bm{B}^{[j,k]}$ are not unique. By multiplying $\bm{B}^{[j,k]}$ with an orthogonal matrix $\bm{U}^{[j,k]}\in\mathbb{R}^{m_{j,k}\times m_{j,k}}$ on the left, we obtain a new $\mathcal{G}$-framelet system $\mathsf{GIF}_{J}(\mathcal{G})$ with the same subspaces $\mathbb{W}_{j,k}$ and preserving the tightness property. Thus, given $\bm{A}^{[j,k]},\bm{B}^{[j,k]}$, to promote the sparsity of $\mathsf{GIF}_{J}(\mathcal{G})$, for each $\mathbb{W}_{j,k}$, we consider solving the following optimization problem:
$\small\begin{gathered}\underset{\bm{U}^{[j,k]}\in\mathcal{M}^{[m_{j,k},m_{j,k}]}}{\arg\min}\ \|\bm{d}^{[j,k]}_{1}(\bm{U}^{[j,k]})^{\top}\|_{1}+\dots+\|\bm{d}^{[j,k]}_{n_{f}}(\bm{U}^{[j,k]})^{\top}\|_{1},\\ \end{gathered}$ (25)
@@
where $\bm{d}^{[0,1]}_{i}=\langle\bm{f}_{i},\bm{\Psi}^{[j,k]}\rangle$, $i\in[n_{f}]$ are the coefficients with respect to $\bm{\Psi}^{[j,k]}$ for $\bm{f}_{i}$, $i\in[n_{f}]$, computed as in Algorithm 3 and $\|\cdot\|_{1}$ is the $l_{1}$ norm. Note that $\bm{d}^{[j,k]}_{1}(\bm{U}^{[j,k]})^{\top}$ are the coefficients in $\mathbb{W}_{j,k}$ with $\bm{B}^{[j,k]}$ replaced by $\bm{U}^{[j,k]}\bm{B}^{[j,k]}$. Therefore, the optimization problem in (25) aims at learning the sparsest basis/frame for a fixed $\mathbb{W}_{j,k}$. To solve (25), we follow [47] and substitute the $l_{1}$ norm $\|\bm{M}\|_{1}$ with a smooth surrogate
$\text{{trace}}\left(\bm{M}^{\top}\tanh(1000\cdot\bm{M})\right),$
@@
where $\tanh$ is applied element-wise. In such a way, similar to (24), the problem in (25) becomes a smooth optimization problem on the Stiefel manifold. Moreover, to reduce the optimization overhead, we do not solve (25) for all $k\in[n_{j}]$ and all $j=0,\ldots,J-1$. In fact, according to the total $l^{2}$ norm of the coefficients, we only solve (25) for certain $j,k$ by choosing among those $\mathbb{W}^{[j,k]}$ the top-$N$ subspaces to optimize. More specifically, we sort the subspaces $\mathbb{W}_{j,k}$ decreasingly according to the initial coefficient quantities $\epsilon^{[j,k]}$:
@@
and solve the optimization problem (25) only for those $j,k$ with respect to the first $N$ largest $\epsilon^{[j,k]}$.
[…]
V Experiments
In this section, we demonstrate the efficiency and effectiveness of our graph framelet systems for the graph signal process through several numerical experiments.
V-A Data Preparation and Implementation Details
@@
    Find a minimal spanning tree of Minnesota traffic graph. Pick a node as the root.
@@
    Generate $n_{d}$ equally-spaced points on interval $[-1,1]$, where $n_{d}$ is the depth of of spanning tree.
@@
    Assign each node with value $T(x_{v_{d}})$ where $v_{d}$ is the depth of the node in the spanning tree and $x_{v_{d}}$ is the $(v_{d}+1)$-th equally-spaced point on $[-1,1]$.
Figure 4: An illustration of Minnesota traffic graph, one of its minimum spanning tree and 2 signal samples from the family $\mathcal{F}$.
Signals from this family are generated by sampling $a_{i}$ from the Guassian distribution and following the evaluation as above (see Figure 4). We generate $50$ signals for training and $5$ signals for testing.
@@
For GraphSS and GraphBior, we adopt the publicly released MATLAB code. In detail, we select “orth” filters and a 3-level decomposition of GraphSS, which results in an orthonormal basis. For GraphBior, we use the same configuration as in the file Biorth_filter bank_demo_2.m, which results in a biorthogonal basis.
@@
The implementation of our method is written in Python [^1] https://github.com/zrgcityu/Graph-Involved-Frame. To construct a partition tree for Minnesota traffic graph, we adopt the Python package scikit-network [^2] https://scikit-network.readthedocs.io and apply a step of post-processing to perform a series of clustering on graphs as described at the beginning of Section II-C. Eventually, we obtain a partition tree with $J=3$, in which each non-leaf node has no less than $15$ children, i.e. $|\mathcal{C}_{j,k}|\geq 15$. This allows choosing large enough $r_{j}$ for $\bm{A}^{[j,k]}$ and thus allows varying the dimension of $\mathbb{V}_{0}$. To obtain a $M$-dimensional $\mathbb{V}_{0}$, we set $r_{J-1}=M$ and the rest $r_{j}$ are set to $1$. As for solving the optimization problem in (24) and the modified version of the optimization problem in (25), we adopt the Python package pymanopt [^3] https://pymanopt.org/[49] and apply conjugate gradient method on Riemannian manifolds. The configuration is set to default and a summary of time consumed is presented in Table I. Figures of graphs are generated using GSPBOX***https://epfl-lts2.github.io/gspbox-html/ [50].
TABLE I: Time consumed on solving Riemannian optimization problems.
Variant Time (Unit: second)
[…]
V-B Non-linear Approximation
[…]
V-B1 Comparison among Variants
@@
Here, we compare the results among variants with different dimensions of $\mathbb{V}_{0}$. The NLA is performed on one of 5 test signals. As shown in Figure 5, when the dimension is $4$, the learned $\mathbb{V}_{0}$ best fit the family $\mathfrak{F}$ and produces the sparsest representation for the test signal, indicating a low-dimensional structure of $\mathfrak{F}$ with a dimension larger than 1. On the other hand, for the cases of dimensions 8 and 12, the results indicate a surplus in defining the size of $\mathbb{V}_{0}$, which distributes the energy of signal over an excessive set and thus results in a denser representation. In contrast, the one-dimension case generally has the densest representation since a one-dimensional $\mathbb{V}_{0}$ is insufficient in fitting the low-dimensional structure of $\mathfrak{F}$.
Figure 5: Relative Error of 4 variants.
[…]
@@
Figure 6: Relative Error of 8 variants.
V-B2 Comparison with Baselines
[…]
Figure 7: Average Relative Error of All Methods.
V-C Denoising
[…]
For each test signal, we add Gaussian noise with $\sigma=1/16,1/8,1/4$ and $1/2$. Then, the coefficients are thresholded with a threshold value set as $3\sigma$. We report the signal-to-noise ratio (SNR). The results are shown in Table II.
Our method surpasses the baselines by a large margin. A sparser representation is better for denoising since more coefficients are dominated by noise, and thus thresholding such coefficients results in less loss of the origin signal. Moreover, we mention in the introduction that frames have the potential to perform better since the energy of noise is distributed over more atoms, resulting in smaller coefficients. With the current setting, as shown in the last column of Table II, this happens when the noise is large enough. In other cases, their performance is very close to the bases.
TABLE II: Average SNR. Best in bold. Unit: dB.

### v3

@@ 4.2 Learning the High-pass Filters
Once the filters $\bm{A}^{[j,k]}$ are determined, in view of the condition in (16), we can obtain each $\bm{B}^{[j,k]}$ by finding a basis or a frame of the orthogonal complement of the rows of $\bm{A}^{[j,k]}$. For the former case, we can apply singular value decomposition (SVD). For the latter case, we can apply the method of constructing undecimated framelet systems in [73, Theorem 2] on a set of orthonormal eigenvectors and select three filters $\bm{\alpha},\bm{\beta}_{1},\bm{\beta}_{2}$ as in [65, Section 4.1]. Here, we choose not to describe the details for brevity and assume that such initial high-pass filters $\bm{B}^{[j,k]}$ are obtained either through the SVD technique or the undecimated framelet approach. We would like to emphasize that any method for constructing frames on the span of a set of orthonormal vectors is eligible.
@@
It is obvious that the high-pass filters $\bm{B}^{[j,k]}$ are not unique. By multiplying $\bm{B}^{[j,k]}$ with an orthogonal matrix $\bm{U}^{[j,k]}\in\mathbb{R}^{m_{j,k}\times m_{j,k}}$ on the left, we obtain a new $\mathcal{G}$-framelet system $\mathsf{GIF}_{J}(\mathcal{G})$ with the same subspaces $\mathbb{W}_{j,k}$ and preserving the tightness property. Thus, given $\bm{A}^{[j,k]},\bm{B}^{[j,k]}$, to promote the sparsity of $\mathsf{GIF}_{J}(\mathcal{G})$, for each $\mathbb{W}_{j,k}$, we consider solving the following problem:
[…]
@@
where $\bm{d}^{[0,1]}_{i}=\langle\bm{f}_{i},\bm{\Psi}^{[j,k]}\rangle$, $i\in[n_{f}]$ are the coefficients with respect to $\bm{\Psi}^{[j,k]}$ for $\bm{f}_{i}$, $i\in[n_{f}]$, computed as in Algorithm 3 and $\|\cdot\|_{1}$ is the $l_{1}$ norm. Note that $\bm{d}^{[j,k]}_{1}(\bm{U}^{[j,k]})^{\top}$ are the coefficients in $\mathbb{W}_{j,k}$ with $\bm{B}^{[j,k]}$ replaced by $\bm{U}^{[j,k]}\bm{B}^{[j,k]}$. Therefore, the problem in (24) aims at learning the sparsest basis/frame for a fixed $\mathbb{W}_{j,k}$. To solve (24), we follow [62] and substitute the $l_{1}$ norm $\|\bm{M}\|_{1}$ with a smooth surrogate
$\operatorname{Tr}\left(\bm{M}^{\top}\tanh(1000\cdot\bm{M})\right),$
@@
where $\tanh$ is applied element-wise. In such a way, similar to (23), the problem in (24) becomes a smooth objective on the Stiefel manifold. Moreover, to reduce the optimization overhead, we do not solve (24) for all $k\in[n_{j}]$ and all $j=0,\ldots,J-1$. In fact, according to the total $l^{2}$ norm of the coefficients, we only solve (24) for certain $j,k$ by choosing among those $\mathbb{W}^{[j,k]}$ the top-$N$ subspaces to optimize. More specifically, we sort the subspaces $\mathbb{W}_{j,k}$ decreasingly according to the initial coefficient quantities $\epsilon^{[j,k]}$:
@@
and solve the problem (24) only for those $j,k$ with respect to the first $N$ largest $\epsilon^{[j,k]}$.
To facilitate the descriptions in the next section, we use $M$ to denote the dimension of $\mathbb{V}_{0}$ and $\mathsf{GIF}$-I(M)(M) to denote a $\mathcal{G}$-framelet system obtained by solving only problem (23) and constructing $\bm{B}^{[j,k]}$ through the undemicated framelet system in [65, Section 4.1]. Similarly, we use $\mathsf{GIF}$-II(M,N)(M,N) to denote a $\mathcal{G}$-framelet system obtained by solving both problem (23) and the modified version of problem (24) on the top-$N$ spaces with respect to the first $N$ largest $\epsilon^{[j,k]}$. When all $\bm{B}^{[j,k]}$ are constructed as bases (from SVD), we denote them slightly differently as $\mathsf{GIB}$-I(M)(M) and $\mathsf{GIB}$-II(M,N)(M,N), respectively.
We conclude this section by some further remarks on the differences between our method and two closely related works [56, 45].
Remark 4.5.
If we replace the first $k$ Laplacian eigenvectors and apply the same procedure in [56], the output will be very similar to ours despite the former being restricted to a basis. Thus, our system seems to be just replacing $k$ Laplacian eigenvectors with $k$ arbitrary orthonormal vectors. In fact, there are some important differences. In [56], the wavelets are constructed on the partition tree in top-down manner, in which the number $k$ affects the shape of the partition tree. In contrast, our method starts from a given partition tree and assigns parameters to each node in the partition in a bottom-up manner. We can adjust the size of parameters so that the number of scaling functions $k$ and the shape of the partition tree are decoupled to a certain extent. This is important since the shape of the partition tree determines the sparseness of the framelets, and we do not want the number $k$ to affect the sparseness. Once the sizes of the parameters are fixed, as shown in Section 4.1, we can determine the values of the parameters in a top-down manner similar to [56] in some special cases.
Remark 4.6.
[45] proposes a framework to construct Parseval frames on directed acyclic graphs. Since the partition tree is an instance of directed acyclic graphs (DAGs), we can essentially produce the same framelets and framelet transforms using the systems in [45] (with slight modifications). However, the notation and data structures adopted in [45] are relatively complicated. In detail, a DAG must be converted into a direct acyclic graph with joint points (DAGJP). A wavelet transform on hierarchical graphs (WTHG), which is a computational diagram depicting the procedure of computing wavelet coefficients, is further obtained by converting the DAGJP. In short, it does not provide comparably simple and concise notations as in classical 1-D framelets (see [21, Chapter 1]) due to its generality. In contrast, when confined to partition trees, our work provides notations and transforms that are more concise and simpler without losing much of the generality. On the other hand, optimizations for unitary transform matrices in [45] are conducted layer-by-layer to obtain sparseness. This is different from the procedure above.
5 Graph Signal Processing
In this section, we demonstrate the efficiency and effectiveness of our graph framelet systems on graph signal processing through several numerical experiments.
5.1 Data Preparation and Implementation Details
@@
    Find a minimal spanning tree of the Minnesota traffic graph. Pick a node as the root.
@@
    Generate $n_{d}$ equally-spaced points on interval [1,1][-1,1], where $n_{d}$ is the depth of of spanning tree.
@@
    Assign each node with value $T(x_{v_{d}})$ where $v_{d}$ is the depth of the node in the spanning tree and $x_{v_{d}}$ is the $(v_{d}+1)$-th equally-spaced point on [1,1][-1,1].
(a) Minnesota traffic
(b) Spanning tree
(c) Sample 1
(d) Sample 2
Figure 3: An illustration of (a) the Minnesota traffic graph, (b) one of its minimum spanning trees, and (c-d) 2 signal samples from the family $\mathfrak{F}$.
Signals from this family are generated by sampling $a_{i}$ from the Gaussian distribution and following the evaluation as above (see Figure 3). We generated $50$ signals for training and $5$ signals for testing.
@@
For GraphSS and GraphBior, we adopted the publicly released MATLAB code. In detail, we selected “orth” filters and a 3-level decomposition of GraphSS, which resulted in an orthonormal basis. For GraphBior, we used the same configuration as in the file Biorth_filter bank_demo_2.m, which resulted in a biorthogonal basis.
@@
Our method was implemented with Python [^1] https://github.com/zrgcityu/Graph-Involved-Frame. To construct a partition tree for the Minnesota traffic graph, we adopted the Python package scikit-network [^2] https://scikit-network.readthedocs.io and applied a step of post-processing to perform a series of clustering on graphs as described at the beginning of Section 2.3. Eventually, we obtained a partition tree with $J=3$, in which each non-leaf node had no less than $15$ children, i.e. $|\mathcal{C}_{j,k}|\geq 15$. This allows choosing large enough $r_{j}$ for $\bm{A}^{[j,k]}$ and thus allows varying the dimension of $\mathbb{V}_{0}$. To obtain a $M$-dimensional $\mathbb{V}_{0}$, we set $r_{J-1}=M$ and the rest $r_{j}$ were set to $1$. As for solving the problem in (23) and the modified version of the problem in (24), we adopted the Python package pymanopt [^3] https://pymanopt.org/[60] and applied the conjugate gradient method on Riemannian manifolds. The configuration was set to default, and a summary of the time consumed is presented in Table 1. Figures of graphs were generated using GSPBOX [^1] https://epfl-lts2.github.io/gspbox-html/ [51].
Table 1: Time consumed on Riemannian optimizations. Unit: second.
Variant Time Variant Time Variant Time
$\mathsf{GIB}$-I(1)(1) 1002.03 $\mathsf{GIB}$-II(1,20)(1,20) 1158.92 $\mathsf{GIB}$-II(1,100)(1,100) 1772.85
$\mathsf{GIB}$-I(4)(4) 1001.59 $\mathsf{GIB}$-II(4,20)(4,20) 1147.13 $\mathsf{GIB}$-II(4,100)(4,100) 1629.42
$\mathsf{GIB}$-I(8)(8) 1002.45 $\mathsf{GIB}$-II(1,50)(1,50) 1388.68 $\mathsf{GIF}$-II(1,20)(1,20) 1366.11
$\mathsf{GIB}$-I(12)(12) 1002.44 $\mathsf{GIB}$-II(4,50)(4,50) 1350.72 $\mathsf{GIF}$-II(4,20)(4,20) 1307.43
5.2 Non-linear Approximation
In this subsection, we choose only bases constructed from our method, e.g., $\mathsf{GIB}$-I(4)(4). To conduct non-linear approximation (NLA) on a signal $\bm{f}$ using top-$N$ coefficients of a dictionary $\mathcal{D}:=\{\bm{g}_{1},\bm{g}_{2},\dots,\bm{g}_{n}\}\subset\mathbb{R}^{n}$, we sorted the coefficients $c_{i}:=\langle\bm{f},\bm{g}_{i}\rangle$ decreasingly according to $|c_{i}|$ and kept the top-$N$ coefficients. The remaining ones were set to 0. Then a signal $\tilde{\bm{f}}$ was reconstructed according to these modified coefficients using the proper inverse operator. We report the relative error, i.e., $\|\bm{f}-\tilde{\bm{f}}\|/\|\bm{f}\|$.
5.2.1 Comparison among Variants
@@
Here, we compare the results among variants with different dimensions of $\mathbb{V}_{0}$. The NLA is performed on one of 5 test signals. As shown in Figure 4 (a), when the dimension is $4$, the learned $\mathbb{V}_{0}$ best fit the family $\mathfrak{F}$ and produces the sparsest representation for the test signal, indicating a low-dimensional structure of $\mathfrak{F}$ with a dimension larger than 1. On the other hand, for the cases of dimensions 8 and 12, the results indicate a surplus in defining the size of $\mathbb{V}_{0}$, which distributes the energy of the signal over an excessive set and thus results in a denser representation. In contrast, the one-dimensional case generally has the densest representation since a one-dimensional $\mathbb{V}_{0}$ is insufficient in fitting the low-dimensional structure of $\mathfrak{F}$.
Figure 4: (a) Relative error of 4 variants of $\mathsf{GIB}$-I. (b) Relative error of 8 variants of $\mathsf{GIB}$-I and $\mathsf{GIB}$-II. (c) Average relative error of all methods.
Given the observation above, we further compare the two extreme cases, i.e. $\mathsf{GIB}$-I(1)(1) and $\mathsf{GIB}$-I(4)(4), for which optimize the filters $\bm{B}^{[j,k]}$ as well. The NLA is performed on the same test signal. As shown in Figure 4 (b), optimizing $\bm{B}^{[j,k]}$ gives sparser representations. For $\mathsf{GIB}$-I(4)(4), the results of optimizing top-20, 50, and 100 subspaces are indistinguishable since the four-dimensional $\mathbb{V}_{0}$ has conserved most of the energy and thus it requires only a small portion of subspaces to optimize.
On the contrary, for $\mathsf{GIB}$-I(1)(1), there is a great portion of energy distributed outside of the one-dimensional $\mathbb{V}_{0}$. Thus, it requires more subspaces to optimize. Only when 100 subspaces are optimized, i.e., $\mathsf{GIB}$-I(1,100)(1,100), does the one-dimensional case achieve a performance comparable with the four-dimensional case, i.e., $\mathsf{GIB}$-II(4,20)(4,20).
@@
5.2.2 Comparison with Baselines
Based on the aforementioned results, we select $\mathsf{GIB}$-II(1,20)(1,20) and $\mathsf{GIB}$-II(4,20)(4,20) to compare with the baselines. We report the average relative error of the 5 test signals. As shown in Figure 4 (c), $\mathsf{GIB}$-II(4,20)(4,20) evidently outperforms all other methods.
5.3 Denoising
In this subsection, we include two more variants $\mathsf{GIF}$-II(1,20)(1,20) and $\mathsf{GIF}$-II(4,20)(4,20) to compare with the baselines.
For each test signal, we added Gaussian noise with $\sigma=1/16,1/8,1/4$ and $1/2$. Then, the coefficients were thresholded with a threshold value set as $3\sigma$. We report the signal-to-noise ratio (SNR). The results are shown in Table 2.
Our method surpasses the baselines by a large margin. A sparser representation is better for denoising since more coefficients are dominated by noise, and thus thresholding such coefficients results in less loss of the original signal. Moreover, we mentioned in the introduction that frames have the potential to perform better since the energy of noise is distributed over more atoms, resulting in smaller coefficients. With the current setting, as shown in the last column of Table 2, this happens when the noise is large enough. In other cases, their performance is very close to the bases.
Table 2: Average SNR. Best in bold. Unit: dB.
@@ 5.3 Denoising
$\mathsf{GIB}$-II(1,20)(1,20) 17.32 12.37 9.49 7.58
$\mathsf{GIB}$-II(4,20)(4,20) 25.31 20.15 14.67 9.62
$\mathsf{GIF}$-II(1,20)(1,20) 16.6 12.16 9.6 7.91
$\mathsf{GIF}$-II(4,20)(4,20) 25.24 20.03 14.67 10.32
6 Node Classification
6.1 Graph Framelets as Features
As mentioned in Section 1.1, we are interested in generating graph framelets to encode structural information other than the adjacency matrices. As the adjacency matrices are simply the connections between nodes and the neighbors that are $1$-hop away, we instead look at structural information that is beyond $1$-hop. To do so, we first introduce the induced two-hop graph. In detail, an induced two-hop graph $\mathcal{G}_{\text{2hop}}:=(\mathcal{V}^{\text{2hop}},\mathcal{E}^{\text{2hop}},\bm{W}^{\text{2hop}})$ of a $n$-node graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\bm{W})$ (assuming it is an undirected and unweighted graph) is defined as $\mathcal{V}^{\text{2hop}}=\mathcal{V}$, $(v_{i},v_{j})\in\mathcal{E^{\text{2hop}}}$ if the shortest path distance between $v_{i},v_{j}$ in graph $\mathcal{G}$ is $2$ for $i,j\in[n]$, and
$\bm{W}^{\text{2hop}}_{i,j}=\begin{cases}1&\text{if $(v_{i},v_{j})\in\mathcal{E_{\text{2hop}}}$,}\\ 0&\text{otherwise,}\end{cases}$
for $i,j\in[n]$. Based on the induced two-hop graph, we have the following definition, which is almost identical to Definition 2.16.
Definition 6.1.
Let $\mathscr{T}_{J}(\mathcal{G}_{\mathrm{2hop}})$ be a multi-graph partition tree associated with an induced $2$-hop graph $\mathcal{G}_{\mathrm{2hop}}$ from $\mathcal{G}$. Let $\bm{A}^{[j,k]}$, $\bm{B}^{[j,k]}\,|\,j=0,\ldots,J-1;k\in[n_{j}]\}$ be the filter matrices satisfying the assumptions of Theorem 2.8. Then for $J_{0}=0,\ldots,J-1$, the $\mathcal{V}^{\mathrm{2hop}}$-framelet system $\mathcal{F}_{J_{0}}^{J}(\mathcal{V}^{\mathrm{2hop}})$ as in (13) is a tight frame for $\mathbb{R}^{n}$, which we call a two-hop framelet (THF) system. In such a case, we simply denote $\mathcal{F}_{J_{0}}^{J}(\mathcal{G}_{\mathrm{2hop}}):=\mathcal{F}_{J_{0}}^{J}(\mathcal{V}^{\mathrm{2hop}})$. In particular, we define for the special case $J_{0}=0$, the THF system $\mathcal{F}_{0}^{J}(\mathcal{G}_{\mathrm{2hop}})$ as
(25) $\mathsf{THF}_{J}(\mathcal{G_{\mathrm{2hop}}}):=\mathcal{F}_{0}^{J}(\mathcal{G}_{\mathrm{2hop}})=\{\bm{\Phi}_{0},\bm{\Psi}_{0},\cdots,\bm{\Psi}_{J-1}\}.$
For simplicity, let $\mathbb{V}_{0}$ be one-dimensional and its basis vector be the constant unit-norm vector. To do so, we let each $\bm{A}^{[j,k]}$ be the constant unit-norm vector of its relevant size. As each $\bm{A}^{[j,k]}$ is decided, according to Lemma 2.1, each $\bm{B}^{[j,k]}$ should be a tight frame of the vector space orthogonal to $\mathrm{span}(\bm{A}^{[j,k]})$. Let $\tilde{\bm{B}}^{[j,k]}$ be the remaining $|\mathcal{C}_{j,k}|-1$ orthonormal vectors in the complement of $\mathrm{span}(\bm{A}^{[j,k]})$, which can be computed using singular value decomposition. We construct a tight frame $\bm{F}^{[j,k]}$ in $\mathbb{R}^{|\mathcal{C}_{j,k}|-1}$ as the coefficient matrix for remaining orthonormal vectors. Then the $\bm{B}^{[j,k]}:=\bm{F}^{[j,k]}\tilde{\bm{B}}^{[j,k]}$ will be our final $\bm{B}$ filters. In details, $\bm{F}^{[j,k]}$ is obtained via frame completion [29] given $3$ vectors of size $|\mathcal{C}_{j,k}|-1$ sampled from Gaussian distribution. In this way, the size of $\bm{F}^{[j,k]}$ is $3(|\mathcal{C}_{j,k}|-1)\times(|\mathcal{C}_{j,k}|-1)$. Frame completion [29] provides the advantage of adjusting the number of framelets without having zero vectors.
Intuitively, each framelet in the two-hop framelet system encodes certain structural information. Moreover, the redundancy in frames results in various framelets being chosen. However, using all framelets will be excessive. Thus, it is necessary to filter the framelets and choose only a certain portion. To do so, we adopt a heuristic approach, in which we measure the variance of each framelet with respect to the normalized Laplacian of the original graph (see (1)), and then sort the variance to produce a “spectral display” of the framelets. The general idea is to integrate two-hop structural information from the two-hop framelets and one-hop structural information from the original graph. Given the spectral display, we can identify the parts of the framelets that provide the most varying patterns. These varying patterns are the features that we expect to be informative and beneficial for node classification. The whole procedure will be presented and discussed in detail in the experiments.
6.2 Experiments
To measure the extent of heterophily, we use the following quantity
$H=\frac{\left|\{(u,v)\in\mathcal{E}|,y_{u}=y_{v}\}\right|}{|\mathcal{E}|}$
where $y_{u},y_{v}$ are the labels of $u,v$. A smaller $H$ indicates a smaller portion of edges that connect nodes from the same classes, and thus a larger heterophily. We conducted experiments on three graphs, amazon-ratings, minesweeper, and tolokers, from [52]. According to the measure above, amazon-ratings possesses relatively larger heterophily, while the other two graphs have small heterophily and could be considered as homophilous graphs. However, as argued in [52], the measure above has certain drawbacks and has high heterophily using other measures. Our proposal of including the graphs minesweeper and tolokers is to further demonstrate the effectiveness of our method in deriving rich structural information, not necessarily restricted to heterophilous graphs.
In forming the partition tree, we controlled that each $|\mathcal{C}_{j,k}|\leq 16$. Then the $\mathsf{THF}_{J}(\mathcal{G}_{\mathrm{2hop}})$ was formed as described in Def. 6.1. Each framelet was further normalized to have unit norm and sorted non-increasingly according to $\bm{f}\tilde{\bm{L}}\bm{f}^{\top}$. The figures in Figure 5 show the overall distribution of variance of the framelets. These are the spectral displays we mentioned in the previous section. We can see that the first and final parts of the diagrams show drastic changes in variance. We considered these parts to have rich patterns and structural information due to the changes in variance. As a result, these framelets were selected as additional features to be input into neural networks. The implementations [^2] https://github.com/zrgcityu/Framelets-as-features were adapted to handle large graphs using sparse matrices.
Figure 5: Distributions of framelet variance. (a): 70299 framelets on amazon-ratings; (b): 29899 framelets on minesweeper; (c): 34185 framelets on minesweeper. Sorted non-decreasingly according to $\bm{f}\tilde{\bm{L}}\bm{f}^{\top}$.
To combine with GNNs, we selected certain framelets with either the lowest or the highest variance in the spectral displays and concatenated them with the original node feature matrices. Our results were obtained based on the best GNNs reported in [52]. In detail, for amazon-ratings, based on a $2$-layer SAGE, we tested framelets with the lowest variance, and the number of framelets takes values in the set $\{1800,1900,2000,2100,2200\}$; for minsweeper and tolokers, based on a $5$-layer GAT-sep, we tested framelets with the highest variance, and the number of framelets takes values in the set $\{100,200,300,400,500\}$. The final results were chosen according to the performance on the validation sets. We adopted the publicly released code from [52] for the GNNs experiments. The results are shown in Table 3, in which the results other than ours are cited from [52]. We can see that for all datasets, combined with our selected framelets as additional features, the GNNs were able to gain a certain amount of improvement and achieve the best results compared with all other methods. We further tested the stability in terms of randomness in generating framelets. As mentioned in Section 6.1, the generation of framelets involves random sampling. Figure 6 shows that for four different random seeds, the framelets resulted in very similar performance, all of which outperformed the second-best result in Table 3.
Table 3: Dataset statistics and classification results. Metric: amazon-ratings, accuracy; minesweeper, tolokers, ROC AUC. Best with bold, second-best with underline.
amazon-ratings minesweeper tolokers
Node 24492 10000 11758
Feature 300 7 10
Edge 93050 39402 519000
Class 5 2 2
$H$ 0.38 0.68 0.59
ResNet [26] 45.90 $\pm$ 0.52 50.89 $\pm$ 1.39 72.95 $\pm$ 1.06
ResNet+SGC 50.66 $\pm$ 0.48 70.88 $\pm$ 0.90 80.70 $\pm$ 0.97
ResNet+adj 51.83 $\pm$ 0.57 50.42 $\pm$ 0.83 78.78 $\pm$ 1.11
H2GCN [76] 36.47 $\pm$ 0.23 89.71 $\pm$ 0.31 73.35 $\pm$ 1.01
CPGNN [75] 39.79 $\pm$ 0.77 52.03 $\pm$ 5.46 73.36 $\pm$ 1.01
GPR-GNN [5] 44.88 $\pm$ 0.34 86.24 $\pm$ 0.61 72.94 $\pm$ 0.97
FSGNN [42] 52.74 $\pm$ 0.83 90.08 $\pm$ 0.70 82.76 $\pm$ 0.61
GloGNN [39] 36.89 $\pm$ 0.14 51.08 $\pm$ 1.23 73.39 $\pm$ 1.17
FAGCN [39] 44.12 $\pm$ 0.30 88.17 $\pm$ 0.73 77.75 $\pm$ 1.05
GBK-GNN [14] 45.98 $\pm$ 0.71 90.85 $\pm$ 0.58 81.01 $\pm$ 0.67
JacobiConv [64] 43.55 $\pm$ 0.48 89.66 $\pm$ 0.40 68.66 $\pm$ 0.65
GCN [30] 48.70 $\pm$ 0.63 89.75 $\pm$ 0.52 83.64 $\pm$ 0.67
SAGE [19] 53.63 $\pm$ 0.39 93.51 $\pm$ 0.57 82.43 $\pm$ 0.44
GAT [63] 49.09 $\pm$ 0.63 92.01 $\pm$ 0.68 83.70 $\pm$ 0.47
GAT-sep 52.70 $\pm$ 0.62 93.91 $\pm$ 0.35 83.78 $\pm$ 0.43
GT [57] 51.17 $\pm$ 0.66 91.85 $\pm$ 0.76 83.23 $\pm$ 0.64
GT-sep 52.18 $\pm$ 0.80 92.29 $\pm$ 0.47 82.52 $\pm$ 0.92
Ours 54.48 $\pm$ 0.27 94.24 $\pm$ 0.44 84.38 $\pm$ 0.57
(a) top l
(b) top r
(c) bottom l
(d) bottom r